OpenStudy (anonymous):

how many five digit zip code numbers are possible if the first digit cannot be a four and adjacent digits cannot be the same?

OpenStudy (anonymous):

the first digit cant be 4, so there are 9 choices the second choice can be a 4, but cannot be the same choice as the first digit . so there are 9 choices for the second digit. for the third digit there are 8 choice, etc so you get 9X9x8x7

OpenStudy (anonymous):

9x9x8x7x6 i mean

OpenStudy (anonymous):

Number theory/discrete mathematics? S = Set of all 5 digit numbers. D = Set of digits 0-9. Know that D^x = "D cross D cross D.... x times) \[S = D^5\]\[n(S) = n(D^5) = 100000\] F = Set of all 5 digit numbers with 4 as the first digit. \[n(F) = n(D^4) = 10000\] A = Set of all 5 digit numbers with 2 or more identical adjacent numbers. \[A = n(D^4) = 10000\](Pretend one of the 4 digits is actually two identical digits next to each other) We're looking for: \[n(S) - n(F \cup A)\] \[n(F \cup A) = n(F) + n(A) - n(F intersect A)\] (sorry, there is no upside-down U... wtf) n(F intersect A) = n(D^3) = 1000 \[ n(F \cup A) = 10000 + 10000 - 1000 = 19000 \]\[n(S) - n(F \cup A) = 100000 - 19000 = 81000\] Your answer is 81,000 possibilities.

OpenStudy (anonymous):

dave thats wrong

OpenStudy (anonymous):

oh crap...

OpenStudy (anonymous):

you did the complement

OpenStudy (anonymous):

I learned the hard way in my discrete mathematics course that a problem that looks extremely simple can turn out to be the most pain-in-the-retricebullsh..t that you never saw coming, and especially if you have to provide a proof..

OpenStudy (anonymous):

yeah i did it too fast , one sec

OpenStudy (anonymous):

i guess i answered a different question

OpenStudy (anonymous):

how did you get ... F intersect A , and how did you find A , you said two adjacent numbers is like 00 123 , etc, but there could be more

OpenStudy (anonymous):

, ok so you grouped it as (00) 1 2 3 , and then you have 10 ^4 possibilities, beause you can have (00) , 10 ways and then times what exactly

OpenStudy (anonymous):

The only thing I'm not positive of in my answer is my evaluation of n(A), so feel free to try and criticize/correct that if you find it is wrong. I'm looking at it now.

OpenStudy (anonymous):

yes Im not so sure with your reasoning there

OpenStudy (anonymous):

i think you overcounted

OpenStudy (anonymous):

Try not to spam the answers, use the chat instead...

OpenStudy (anonymous):

lets look at the complement, n(A) = 1 - n(A')

OpenStudy (anonymous):

woops, im doing probability. n(A) = universe - n(A') , n(A') = 10*9*8*7*6 ,

OpenStudy (anonymous):

universe = n(D^5) = 10^5, so n(A) = 10^5 - 10*9*8*7*6

OpenStudy (anonymous):

you sound like somebody who is either extremely rusty in the topic, or only knows a little.. I'm guessing rusty? you don't subtract cardinalities from the Universe.. the universe is a set and cardinality is an integer

OpenStudy (anonymous):

regardless, i think you may be going in the right direction with factorials

OpenStudy (anonymous):

the universe is S . , so n(A) = n(S) - n(A')

OpenStudy (anonymous):

there ya go

OpenStudy (anonymous):

wow, that was so important

OpenStudy (anonymous):

omg, the whole problem hinged on that

OpenStudy (anonymous):

but let me give you my reasoning, do you agree with n(A')

OpenStudy (anonymous):

the cardinality of having no adjacent numbers. the first choice is 10 digits, then the next can only by 9, then 8 , etc

OpenStudy (anonymous):

no because this reasoning doesn't allow a number such as 12121, which does not have any same adjacent numbers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hm okay how about 10*9*9*9*9

OpenStudy (anonymous):

for n(A') ?

OpenStudy (anonymous):

i think that may be the solution right there.. sooo 65610

OpenStudy (anonymous):

yeah, and that would make sense.. now we just have to find the intersection between that and F

OpenStudy (anonymous):

which you could do by skipping the whole F deal, and just making it 9^5 i think?

OpenStudy (anonymous):

ah no it's wrong

OpenStudy (anonymous):

i agree thats 10*9*9*9*9 for n(A')

OpenStudy (anonymous):

so n(A) = n(S) - n(A') ,

OpenStudy (anonymous):

9^5 = 59, 049

OpenStudy (anonymous):

ah maybe it's right.. haha aghhhh

OpenStudy (anonymous):

9 choices for first number (can't be 4) 9 choices for second number (can't be first number) 9 choices for third number (can't be second number) 9 choices for fourth number (can't be third number) 9 choices for fifth number (can't be fourth number) I also verified it with a short program.

OpenStudy (anonymous):

right, thats logical

OpenStudy (anonymous):

^Thank you for doing what I was too lazy to do...

OpenStudy (anonymous):

my first answer assumed no repeating for some reason, woops

OpenStudy (anonymous):

thats by multiplication principle

OpenStudy (anonymous):

dave, you botched this one pretty bad

OpenStudy (anonymous):

and then you tried to insult me. haha

OpenStudy (anonymous):

im not even sure about n(F intersect A)

OpenStudy (anonymous):

you got 10^3 for that , hmmmm

OpenStudy (anonymous):

i asked you about your past experience.. not trying to insult you. this isn't a social playground, we're just trying our best to help somebody get the correct answer here

OpenStudy (anonymous):

the person is long gone

OpenStudy (anonymous):

who asked the question

OpenStudy (anonymous):

anyways, how about finding n(A) now ?

OpenStudy (anonymous):

n(A) = 34390

OpenStudy (anonymous):

directly finding n(A) is difficult,

OpenStudy (anonymous):

dmancine, how did you verify it with a short program?

OpenStudy (anonymous):

ok to find n(A) directly you split it into cases, find when there are exactly 2 adjacent, then exactly 3 adjacent, ... exactly 5 adjacent. these cases are mutuall exclusive and you can use addition rule

OpenStudy (anonymous):

http://codepad.org/MMK9wzUM